Find the x-intercepts.
y=x^2-4x+4
Thank you.

Answer by mukhopadhyay(490)

x-intercept is found by equating y=0;
y=x^2-4x+4
=> x^2-4x+4 = 0
=> (x-2)^2 = 0
x-intercept is (2,0). There is only one x-intercept here.